OpenStudy (anonymous):

Differentiate. F(y) = 1/(y^2 −3/y^4)(y + 5y^3)

OpenStudy (anonymous):

F(y) = (1/y^2 −3/y^4)(y + 5y^3)

OpenStudy (anonymous):

use product rule?

OpenStudy (anonymous):

is this the correct equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

F(y) = (1/y^2 −3/y^4)(y + 5y^3)

OpenStudy (anonymous):

\[F(y)=\frac{ 1 }{ y ^{2} -\frac{ 3 }{ y ^{4} }}(y+5y ^{3})\]

OpenStudy (anonymous):

how does this one look?

OpenStudy (anonymous):

\[\frac{ 1 }{ y ^{2} }-\frac{ 3 }{ y ^{4} }(y+5y ^{3})\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

both fractions r in a bracket together

OpenStudy (anonymous):

\[F(y)=y ^{-2}-3y ^{-4}(y+5y ^{3})\] I would first rewrite it like this and then use the product/chain-rule.

OpenStudy (anonymous):

okkay

OpenStudy (anonymous):

(y^−2−3y^−4)(y+5y3)

OpenStudy (anonymous):

\[F'(y)=-2y ^{-3}-3y ^{-4}(1+15y ^{2})+(y+5y ^{3})(12y ^{-5})\]

OpenStudy (anonymous):

you can combine like terms in order to get a simplified answer

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

you're welcome